Can you solve for N rather than L?
L=(R^2*N^2)/((9*R)+(10*(N/P)))
I'm stumped.
Thanks for your attention

Answer by rothauserc(4718)

L=(R^2*N^2)/((9*R)+(10*(N/P)))
(R^2 * N^2) = L * ((9*R)+(10*(N/P)))
(R^2 * N^2) = L * ((9RP +10N) / P)
(R^2 * N^2) = (9RPL +10NL) / P
P*R^2*N^2 = 9RPL +10NL
N^2 = 9RPL / PR^2 + 10NL / PR^2
N^2 - 10NL / PR^2 = 9RPL / PR^2
(N^2 * PR^2 -10NL) / PR^2 = 9RPL / PR^2
N^2 * PR^2 -10NL = 9RPL
P*N^2*R^2 -10NL = 9RPL
N*(P*N*R^2 -10L) = 9RPL